OpenStudy (anonymous):

What is the value of the expression? 19.2 + (–7 + 3.8) A. 16 B. 24 C. 29 D. 30

OpenStudy (anonymous):

A (16)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how about What is the value of the expression? 19.2 + (���������7 + 3.8) …

OpenStudy (anonymous):

–(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (anonymous):

all that came up was a bunch of "?"

OpenStudy (anonymous):

it is –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (anonymous):

I don't know that one, sorry!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

–(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17